# On The Computational Complexity for Minimizing Aerial Photographs for Full Coverage of a Planar Region

### v1

@@ 1 Sample NN points from the polygon
Limitations and future work
In this work, we showed the inapproximability ratio of $1.165$ for using the minimum side length of $k$ squares to cover the boundary of a simple polygon. We also considered the constrained cases when the center of the squares or circles should lie inside the region or boundary, which in these cases the inapproximability ratio increases to $1.25$ for the square coverage problem.
The current gap between the $1.152$-inapprox. ratio and the efficient $2$-approx. algorithm for the circle coverage problem, and the gap between the $1.165$-inapproximability ratio and the efficient $2.828$-approx. algorithm for the square coverage problem is still large. Same for the gap between the restricted version of the square and circle coverage problem having $1.152$ and $1.25$ in-approximability ratio respectively and $2$-approx. and $2.828$-approx. algorithms respectively. So, there still exist much space for improvements by proving higher inapproximability ratios or by developing approximation algorithms with better guarantees.
@@ Acknowledgements
The author would like to thank the colleagues at Autel Robotics, especially Dr. Kartik Mohta, Dr. Kangli Wang, Dr. Yan Lu, and Keke Zhang, for the help on providing opinions on the aerial robotics context of the problem studied in this paper, and for providing intuitions for this study with the high-quality and high-zoom-in capability gimbal cameras. Prof. Jingjin Yu’s help and brain-storming discussions at the start of this line of work are essential to the results developed in this paper.

### v2

@@ 1 Sample NN points from the polygon with sampling density $\epsilon$
Discussions and future work
In this work, we showed the inapproximability ratio of $1.165$ for using the minimum side length of $k$ squares to cover the boundary of a simple polygon. We also considered the constrained cases when the center of the squares or circles should stay inside the region or boundary, which in these cases the inapproximability ratio increases to $1.25$ for the square coverage problem.
The current gap between the $1.152$-inapprox. ratio and the efficient $2$-optimal algorithm for the circle coverage problem, and the gap between the $1.165$-inapprox. ratio and the efficient $2.828$-optimal algorithm for the square coverage problem are still large. Same for the gap between the restricted version of the square and circle coverage problem having $1.152$ and $1.25$ inapproximability ratio respectively and $2$-optimal and $2.828$-optimal algorithms respectively. So, there still exists much space for improvements by proving higher inapproximability ratios or by developing approximation algorithms with better guarantees.
@@ Acknowledgements
The author would like to thank the colleagues at Autel Robotics, especially Dr. Kartik Mohta, Dr. Kangli Wang, Dr. Yan Lu, and Keke Zhang, for the help on providing opinions on the aerial robotics context of the problem studied in this paper, and for providing intuitions for this study with the high-quality and high-zoom-in-capability gimbal cameras. Prof. Jingjin Yu’s help and brainstorming discussions at the start of this line of work are essential to the results developed in this paper.